# Composing Agents to Minimize Worst-case Risk

### v1

@@ 5 Limitations
A limitation of the BucketedVaR algorithm is that we require the losses of agents to be independent to obtain tight estimates of the value-at-risk. For instance, this is violated when an agent produces a safe trajectory and ends in a state from which the next agent cannot be safe. In such cases, we are forced to rely on the inefficient baseline algorithm. Yet another limitation of our framework is that we may not always have well defined loss functions that quantify risk like in the case of an LLM agent. Here, we may have to rely on LLM-as-a-judge methods or human annotations which can be expensive and noisy.
@@ 6 Conclusion and future directions
In summary, we introduced a framework for choosing optimal agent compositions to minimize worst-case risk criteria that can encode safety, privacy, and fairness desiderata. Our dynamic programming approach efficiently allocates the risk budget across agents, supporting tight approximations under mild independence assumptions. Empirical results on compositional RL benchmarks confirm its effectiveness in identifying optimal paths and quantifying tail behaviors, paving the way for more reliable agentic workflows.
A direction for future work is to incorporate the BucketedVaR algorithm in existing agentic frameworks and evaluate it on LLM-agents. On the theoretical front, an interesting direction is to explore whether we can develop online learning algorithms to minimize worst-case risk. Such an algorithm would have to pick a composition of agents at every turn and minimize long-term regret with respect to the optimal composition.
@@ Appendix A Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
The optimal allocation would ensure that the $(1-\alpha_{i}^{*})$-quantile is the same for all edges, which we’ll denote as $q^{*}$. This means $F_{i}(q^{*})=1-\alpha_{i}^{*}$ for all $i$. Since BucketedVaR searches over discretized buckets, as $d\rightarrow\infty$, it approaches this optimal allocation.
@@ Appendix B Benchmark environments and experimental evaluation setups
This benchmark environment is borrowed from Jothimurugan et al. [17]. The state and action spaces are continuous and involves controlling a point mass to navigate between rooms while avoiding obstacles. A map of the environment is shown is shown in Figure 4. Jothimurugan et al. [17] specify objectives in the form of task graphs where each vertex corresponds to a set of subgoal states and each edge represents a reach-avoid subtask of reaching the target vertex’s subgoal states. The task graph that we consider corresponds to the specification $\phi_{5}$ from Jothimurugan et al. [17] and is also visualized in Figure 4. It consists of 13 vertices arranged in a sequence of 4 diamonds with 4 vertices in each diamond like in Figure 1. The first diamond specifies starting from the light blue square and reaching the room marked with the dark blue square in Figure 4 using either the top path while avoiding the red obstacle or using the bottom path. Similarly, the next diamond specifies the two paths to reach the green square, followed by the pink, and brown squares. Each edge in the task graph is also associated with a real-valued reward function that evaluates to at least zero when a trajectory of the controller satisfies the reach-avoid subtask. This is function of the distance between a trajectory and the obstacles and between the trajectory and the subgoal region. So if the reward is positive then it means that the trajectory stays at least a certain distance away from the obstacles and reaches at most a certain distance to the goal.
@@
In Table 1, since the estimated VaR at all levels for the 16-Rooms benchmark is negative, it implies that with probability at least $1-\alpha$, the reward along all the edges of the path are at least the positive estimated VaR, where $\alpha$ is the risk level.
@@ B.2 Fetch
B.3 MouseNav
As described previously, we adapt the 16-Rooms environment to implement the MouseNav environment from Figure 1. Additionally, we adjust the sizes of the obstacles so that one of the paths has bigger obstacles than the other path.
@@ B.3 MouseNav
This environment was designed as a counterexample—the algorithm from Jothimurugan et al. [17] would be unable to pick a path in the graph since it only maximizes the success probability and in this case both are feasible to complete the reach-avoid tasks with probability 1. On the other hand, optimizing the WoRM objective would give us that the path with the smaller obstacles is safer.
@@ B.4 BoxRelay
The environment and the agent graph details are presented in Section 4. The policies are trained using the PPO [28] CNN-Policy algorithm from StableBaselines3 package with the default hyperparameters for 500,000 iterations for each edge policy. The initial state distribution was again estimated by taking rollouts of the previous edge policy. The loss functions in the agent graph are defined as follows where the vertices are numbered as shown in Figure 2(a).

### v2

@@ 5 Limitations
A limitation of our algorithm is that we require the losses of agents to be independent to obtain the theoretical guarantee regarding the tightness of the value-at-risk estimates. Empirically, however, we observe that it continues to produce tight estimates even without any formal independence guarantees which indicates that we may be able to relax this assumption. Regardless, we emphasize that this independence assumption holds when the losses are metrics of the agent behavior rather than task performance which is the case in all our benchmarks and examples. Another limitation is that we may not always have well defined loss functions like in the case of an LLM agent and we may have to rely on LLM-as-a-judge methods or human annotations which may be expensive and noisy.
@@ 6 Conclusion and future directions
In summary, we introduced a framework for choosing optimal agent compositions to minimize the value-at-risk of losses that encode violations of safety, fairness, and privacy requirements. Our dynamic programming approach efficiently allocates the risk budget across agents, supporting tight approximations under mild independence assumptions. Empirical results on compositional RL benchmarks confirm its effectiveness in identifying optimal paths and quantifying tail behaviors, paving the way for more reliable agentic workflows.
A direction for the future is to incorporate the BucketedVaR algorithm in existing agentic frameworks. It is also possible to speed up the algorithm through a GPU implementation since the computation of the intermediate VaR estimates is highly parallelizable. While we present an efficient algorithm to minimize the value-at-risk in agent graphs, it is also an interesting future direction to extend our work to the estimation of other popular risk measures like the expected shortfall (CVaR) and the class of coherent risk measures [2]. Yet another an interesting direction is to explore whether we can develop online multi-armed bandit style algorithms to minimize risk. Such an algorithm would have to pick a composition of agents at each turn and minimize long-term regret with respect to the optimal composition. Recent works in risk-aware bandits [28] and combinatorial bandits [3] seem to indicate that this is feasible.
@@ Appendix A Proofs of correctness, efficiency, and optimality of the BucketedVaR algorithm
The optimal allocation would ensure that the $(1-\alpha_{i}^{*})$-quantile is the same for all edges, which we denote as $q^{*}$. This means $F_{i}(q^{*})=1-\alpha_{i}^{*}$ for all $i$. Since BucketedVaR searches over discretized buckets, as $d\rightarrow\infty$, it approaches this optimal allocation.
@@ Appendix B Benchmark environments and experimental evaluation setups
This benchmark environment is borrowed from Jothimurugan et al. [18]. The state and action spaces are continuous and involves controlling a point mass to navigate between rooms while avoiding obstacles. A map of the environment is shown is shown in Figure˜5. Jothimurugan et al. [18] specify objectives in the form of task graphs where each vertex corresponds to a set of subgoal states and each edge represents a reach-avoid subtask of reaching the target vertex’s subgoal states. The task graph that we consider corresponds to the specification $\phi_{5}$ from Jothimurugan et al. [18] and is also visualized in Figure˜5. It consists of 13 vertices arranged in a sequence of 4 diamonds with 4 vertices in each diamond like in Figure˜1. The first diamond specifies starting from the light blue square and reaching the room marked with the dark blue square in Figure˜5 using either the top path while avoiding the red obstacle or using the bottom path. Similarly, the next diamond specifies the two paths to reach the green square, followed by the pink, and brown squares. Each edge in the task graph is also associated with a real-valued reward function that evaluates to at least zero when a trajectory of the controller satisfies the reach-avoid subtask. This is function of the distance between a trajectory and the obstacles and between the trajectory and the subgoal region. So if the reward is positive then it means that the trajectory stays at least a certain distance away from the obstacles and reaches at most a certain distance to the goal.
@@
In Table˜1, since the estimated VaR at all levels for the 16-Rooms benchmark is negative, it implies that with probability at least $1-\alpha$, the reward along all the edges of the path are at least the positive estimated VaR, where $\alpha$ is the risk level.
@@ B.2 Fetch
B.3 DroneNav
As described previously, we adapt the 16-Rooms environment to implement the DroneNav environment from Example 1. Additionally, we adjust the sizes of the obstacles so that one of the paths has bigger obstacles than the other path.
@@ B.3 DroneNav
This environment was designed as a counterexample—the algorithm from Jothimurugan et al. [18] would be unable to pick a path in the graph since it only maximizes the success probability and in this case both are feasible to complete the reach-avoid tasks with probability 1. On the other hand, optimizing the RMAG objective would give us that the path with the smaller obstacles is safer.
@@ B.4 BoxRelay
Figure 6: First-person view when carrying box in the BoxRelay environment which also serve as observations.
The environment and the agent graph details are presented in Section˜4. The policies are trained using the PPO [27] CNN-Policy algorithm from StableBaselines3 package with the default hyperparameters for 500,000 iterations for each edge policy. The initial state distribution was again estimated by taking rollouts of the previous edge policy. The loss functions in the agent graph are defined as follows where the vertices are numbered as shown in Figure˜2.
@@ B.5 Experiment computational requirements
Appendix C Experimental results
Table 1: Comparison of BucketedVaR with the optimal baseline algorithm on $10^{4}$ samples. We prefer empirical quantiles that are close to $1-\alpha$ where $\alpha$ is the risk level. 95% Clopper-Pearson CIs are specified for the empirical quantiles computed on a fresh set of $10^{4}$ samples.
Benchmark
Risk
level
Buckets
BucketedVaR
estimate
Baseline
estimate
BucketedVaR
quantile (%)
Baseline
quantile (%)
DroneNav 0.2 5 -0.339 -0.339 79.56[78.75, 80.34] 79.56[78.75, 80.34]
0.1 5 -0.330 -0.330 89.77[89.15, 90.35] 89.77[89.15, 90.35]
0.05 5 -0.328 -0.328 95.01[94.56, 95.42] 95.01[94.56, 95.42]
0.1 100 -0.0135 -0.0135 91.24[90.66, 91.78] 91.24[90.66, 91.78]
0.05 100 -0.0084 -0.0084 95.45[95.02, 95.85] 95.45[95.02, 95.85]
Fetch 0.2 30 0.0314 0.0250 80.45[79.65, 81.22] 79.58[78.77, 80.36]
0.1 30 0.3483 0.3393 90.92[90.33, 91.47] 90.53[89.93, 91.09]
0.05 30 0.3914 0.3881 95.40[94.97, 95.80] 95.28[94.84, 95.68]
BoxRelay 0.2 50 151 148 81.79[81.01, 82.54] 80.56[79.77, 81.33]
0.1 50 176 175 90.90[90.31, 91.45] 90.70[90.11, 91.26]
0.05 50 210 208 95.46[95.03, 95.85] 95.34[94.90, 95.74]
Complete evaluation results of the BucketedVaR algorithm along with comparison with the baseline optimal algorithm are given in Table˜1.